OpenStudy (anonymous):

the product of 2 and a number increased by 7 is -36 Translate then solve need help please

OpenStudy (anonymous):

just read carefully. double a number and add 7. it equals -36

OpenStudy (anonymous):

you want to write an equation here

OpenStudy (saifoo.khan):

Hint: suppose that number be x

OpenStudy (anonymous):

I thought it would be 2x+7=-36 but answer doesnt come out right

OpenStudy (saifoo.khan):

Product = Multiply. Sum = Add

OpenStudy (unklerhaukus):

\[2 \times n + 7 = -36\] \[n=-43/2\]